OpenStudy (anonymous):

Solve the system by substitution. -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (anonymous):

well where are you having problem?

OpenStudy (anonymous):

i can graph the equation but im not sure what they mean but solve with substitution

OpenStudy (anonymous):

okay let's take first equation can you solve it for x?

OpenStudy (anonymous):

do you combine like terms?

OpenStudy (anonymous):

-4.5x - 2y = -4.5x - 2(3.25x-.75) = 12.15 this is what i have so far

OpenStudy (anonymous):

yes you are right.. now as you have equation in terms of x you can solve it -4.5x-2(3.25x-.75=12.15 further simplifying -4.5x-6.5x-1.5=12.15 -11x=12.15+1.5 can you solve this now?

OpenStudy (anonymous):

x=-1.24